OpenStudy (anonymous):

HELP d/690=15 5x+3=23

OpenStudy (anonymous):

to solve for d multiply both sides by 690 so that d = 15 times 690 = then for 5x + 3 = 23 subtract 3 from both sides -3 -3 so 5x = 20 then divide both sides by 5 /5 /5 so x = 20 divided by 5 = (simplify)

OpenStudy (anonymous):

the first one is d=690=115 sorry. so is it 115 times 690?

OpenStudy (anonymous):

*d/690